# {centering} Higgs pole inflation with loop corrections in light of ACT results

### v1

@@
We present the Coleman-Weinberg potential for the inflaton in the pole inflation scenarios such as the Higgs pole inflation and the Peccei-Quinn (PQ) pole inflation. The loop corrections stem from the Standard Model particles and extra singlet scalar fields in the former case, making the quartic coupling for the Higgs inflaton modified by the inflaton-dependent power corrections during inflation. We also obtain similar power corrections to the quartic coupling for the PQ inflaton, depending on the realizations of the PQ symmetry in KSVZ and DFSZ models. We show that the loop corrections shift the spectral index in the pole inflation to a larger value in favor of the ACT results, while being compatible with the bound on the tensor-to-scalar ratio. For a positive one-loop beta function for the inflaton quartic coupling (namely, $b_{1}>0$), a sub-dominant contribution from the two-loop corrections can be accommodated. On the other hand, if the one-loop beta function for the inflaton coupling is negative (namely, $b_{1}<0$), we need sizable contributions from two-loops that are larger than the one-loop corrections due to the ACT results.
@@ 1 Introduction
Considering the renormalization group improved inflaton potential in both Higgs and PQ inflation models, we parametrize the effects of the loop corrections as a running quartic coupling for the inflaton at the two-loop order. Then, we analyze the impacts of the loop corrections on the spectral index and the tensor-to-scalar ratio in the perturbativity regime where the running quartic coupling does not change significantly during inflation. As a result, we discuss the implications of the ACT results for the beta function coefficients for the inflaton quartic coupling and how the couplings of the inflaton to new particles or the content of the models are constrained.
@@ 3.2 Coleman-Weinberg potential for the Higgs inflaton
where $\alpha=\{Z,W,t,h,G\}$ for $Z,W$ gauge bosons, top quark, Higgs and Goldstones, respectively, with $N_{\alpha}=\{3,6,-12,1,3\}$, and $C_{\alpha}=\frac{3}{2}$ for fermions or scalars and $C_{\alpha}=\frac{5}{6}$ for gauge bosons. For large Higgs field values during inflation, we choose the renormalization scale to $\mu=c_{*}h/\sqrt{\Omega}$ with $c_{*}$ being the optimal value of order one in order to minimize the logarithms in the Coleman-Weinberg potential [18]. Then, following the procedure described in the Appendix A, we find that the renormalized effective potential during inflation is expanded for $h^{2}<6$ as
@@
where the running quartic coupling $\lambda_{H}$ becomes field-dependent due to $\mu=c_{*}h/\sqrt{\Omega}$ and we introduced the sextet coupling $\lambda_{6}$ for the Higgs after the renormalization of the CW potential  [^1] See the details for the renormalization of the CW potential at one-loops in the Appendix A.. Henceforth, assuming that the renormalized sextet coupling and the higher order self-couplings for the Higgs are set to zero, we focus on the running Higgs quartic coupling. We note that the Higgs and Goldstone contributions to the CW potential can be suppressed, relative to $W,Z$ boson and top quark contributions, but they are important for deriving the RG equations as shown in the Appendix A.
@@
The running quartic coupling can be nearly zero around the inflation scale even within the SM, depending on the top quark mass, and it is also possible to make the one-loop beta function for the Higgs quartic coupling with new physics contributions to be small around the inflation scale. In this case, we regard the Higgs quartic coupling as being running up to two-loops near the inflation scale $\mu_{I}$, as follows,
@@ 3.3 The case for the PQ inflaton
with the field-dependent quartic coupling $\lambda_{\Phi}(\rho)$ and the sextet coupling ${\bar{\lambda}}_{6}$ introduced after the renormalization of the CW potential. Similarly to the Higgs pole inflation, we set the renormalized sextet coupling and the higher order self-couplings for the PQ field to zero and focus on the running quartic coupling for the PQ field up to two-loops as in eq. (29). In this case, the beta function coefficients depend on $y_{Q},y_{N}.,\lambda_{H\Phi}$ and the SM couplings in KSVZ models and $y_{N},\kappa,\lambda_{1\Phi},\lambda_{2\Phi}$ and the SM couplings in DFSZ models. A small running quartic coupling for the PQ field is maintained due to small couplings of the PQ fields, namely, as far as $y_{Q},y_{N}\lesssim 10^{-3}$ and $\lambda_{H\Phi},\kappa,\lambda_{1\Phi},\lambda_{2\Phi}\lesssim 10^{-5}$ [14, 15].
@@ 4.2 Inflationary observables
[…]
@@
[…]
@@
The observed spectral index is given by $n_{s}=0.9649\pm 0.0044$ from Planck [2], $n_{s}=0.9709\pm 0.0038$ from Planck$+$ACT [3], and $n_{s}=0.9743\pm 0.0034$ from Planck$+$ACT$+$LB [3]. On the other hand, the bound on the tensor-to-scalar ratio from the combined Planck and Keck data is given by $r<0.036$ at $95\%$ CL [6].
@@
Figure 1: Parameter space for $b_{1}/\lambda_{H_{*}}$ vs $b_{2}/\lambda_{H_{*}}$ being consistent with Planck$+$ACT$+$LB data [3]. The range of the spectral index within $1\sigma$ and $2\sigma$ errors are shown in yellow and green. We took $N=50$ on left and $N=60$ on right. The gray region is not consistent with perturbativity and the purple region shows the two-loop dominance.
@@
Figure 3: The spectral index vs the tensor-to-scalar ratio. The range of the spectral index allowed by Planck$+$ACT$+$LB data [3] are shown in yellow and green colors within $1\sigma$ and $2\sigma$ errors, respectively. We chose $b_{1}/\lambda_{H_{*}}=0.02(-0.01)$ and $b_{2}/\lambda_{H_{*}}=0.001(0.005)$ on left and right, respectively.
In Fig. 1, we depict the parameter space for the one-loop and two-loop beta function coefficients, namely, $b_{1}/\lambda_{H_{*}}$ vs $b_{2}/\lambda_{H_{*}}$, that are both positive and are consistent with Planck$+$ACT$+$LB within $1\sigma$ and $2\sigma$ errors [3]. We took $N=50,60$ in the left and right plots, respectively. The gray region is not consistent with perturbativity conditions for loop corrections while the purple region is where the two-loop corrections are larger than the one-loop corrections.
@@
In Fig. 3, we also show the predictions for the spectral index vs the tensor-to-scalar ratio in our model. We chose $b_{1}/\lambda_{H_{*}}=0.02$ and $b_{2}/\lambda_{H_{*}}=0.001$ in the left plot while $b_{1}/\lambda_{H_{*}}=-0.01$ and $b_{2}/\lambda_{H_{*}}=0.005$ in the right plot. For $N=50-60$, the inflationary predictions can be read along the black line bounded by two black bullets. In these examples, for $N=50-60$, the spectral index and the tensor-to-scalar ratio vary by $n_{s}=0.967-0.974$ and $r=0.0050-0.0068$ on left and $n_{s}=0.966-0.973$ and $r=0.0040-0.0057$ on right. As a result, the running quartic coupling at the horizon exit in eq. (58) varies by $\lambda_{H,*}=(1.7-2.3)\times 10^{-11}$ on left and $\lambda_{H,*}=(1.4-2.0)\times 10^{-11}$ on right. The results are in contrast with the case at tree level for $b_{1}=b_{2}=0$: $n_{s}=0.960-0.967$ and $r=0.0033-0.0048$ for $N=50-60$, so the predictions for the spectral index are in a tension with Planck$+$ACT$+$LB at about $2-4\sigma$ level.
@@ 5 Conclusions
We presented the Coleman-Weinberg potential for the inflaton in the pole inflation scenarios through the running quartic coupling for the inflaton. The loop corrections stem from the SM particles and extra singlet scalar fields in the Higgs pole inflation, giving rise to the inflaton-dependent power corrections to the Higgs quartic coupling during inflation, although the non-inflaton fields are much heavier than the inflaton. Similar power corrections are obtained for the PQ pole inflation, due to the couplings of the PQ field in the realizations of the PQ symmetry: the Yukawa couplings of the heavy vector-like quark and the right-handed neutrinos and the portal coupling to the SM Higgs doublet in KSVZ models, and the Yukawa couplings of the right-handed neutrinos and the portal couplings to two Higgs doublets in DSFZ models.
We showed that the loop corrections make the inflationary predictions in the pole inflation deviating from the tree-level results, in favor of the ACT results. For a positive one-loop beta function for the inflaton quartic coupling (namely, $b_{1}>0$), a sub-dominant contribution from the two-loop corrections can be accommodated. On the other hand, if the one-loop beta function for the inflaton coupling is negative (namely, $b_{1}<0$), we need sizable contributions from two-loops that are larger than the one-loop corrections to explain the ACT results. Whether $b_{1}>0$ or $b_{1}<0$, we found that the tensor-to-scalar ratio remains small enough to be compatible with the bound from Planck and Keck, although it gets slightly larger due to the loop corrections.

### v2

@@
We present the Coleman-Weinberg potential for the inflaton in the pole inflation scenarios such as the Higgs pole inflation and the Peccei-Quinn (PQ) pole inflation. The loop corrections stem from the Standard Model particles and extra singlet scalar fields in the former case, making the quartic coupling for the Higgs inflaton modified by the inflaton-dependent power corrections during inflation. We also obtain similar power corrections to the quartic coupling for the PQ inflaton, depending on the realizations of the PQ symmetry in KSVZ and DFSZ models. We show that the loop corrections can shift the spectral index in the pole inflation to a larger value in favor of the ACT results, while being compatible with the bound on the tensor-to-scalar ratio. For a positive one-loop beta function for the inflaton quartic coupling (namely, $b_{1}>0$), a sub-dominant contribution from the two-loop corrections can be accommodated. On the other hand, if the one-loop beta function for the inflaton coupling is negative (namely, $b_{1}<0$), we need sizable contributions from two-loops that are larger than the one-loop corrections due to the ACT results.
@@ 1 Introduction
Considering the renormalization group (RG) improved inflaton potential in both Higgs and PQ inflation models, we parametrize the effects of the loop corrections as a running quartic coupling for the inflaton at the two-loop order. Then, we analyze the impacts of the loop corrections on the spectral index and the tensor-to-scalar ratio in the perturbative regime where the running quartic coupling does not change significantly during inflation. As a result, we discuss the implications of the ACT results for the beta function coefficients for the inflaton quartic coupling and how the couplings of the inflaton to new particles or the content of the models are constrained. There are previous studies on the RG improved inflaton potential in Higgs-like inflation models [17].
@@ 2 Setup
We note that the form of the Jordan-frame Higgs potential in eq. (5) is motivated from the Weyl invariant Lagrangian of the Higgs inflation [16]. In this case, the form of the Higgs potential is related to the form of the non-minimal coupling to gravity in the limit of the non-compact isometry $SO(1,N)$ in the field space for the dilaton $\chi$ and the Higgs doublet $H$ with $N=4$ (or the PQ field with $N=2$). For instance, the Weyl invariant Lagrangian for the Higgs doublet with an approximate $SO(1,4)$ isometry [16] is given by
$\frac{{\cal L}_{J}}{\sqrt{-g_{J}}}$ $=$ $(1+a)\bigg[-\frac{1}{12}(\chi^{2}-2|H|^{2})R-\frac{1}{2}(\partial_{\mu}\chi)^{2}+|D_{\mu}H|^{2}\bigg]$ (7)
$+\frac{1}{2}a({\hat{D}}_{\mu}\chi)^{2}-a|{\hat{D}}_{\mu}H|^{2}-\frac{1}{4}w_{\mu\nu}w^{\mu\nu}-V_{J}(\chi,H)$
where $w_{\mu}$ is the Weyl gauge field, $a$ is a constant parameter, ${\hat{D}}_{\mu}\chi=(\partial_{\mu}-g_{w}w_{\mu})\chi$, ${\hat{D}}_{\mu}H$ contains the covariant derivative containing Weyl gauge field by ${\hat{D}}_{\mu}H=(\partial_{\mu}-g_{w}w_{\mu}+\cdots)H$, and $D_{\mu}H$ is the covariant derivative in the SM. Here, the Jordan frame potential takes
$V_{J}(\chi,H)=\frac{1}{\langle\chi^{4}\rangle}\,f(|H|^{2}/\chi^{2})(\chi^{2}-2|H|^{2})^{2},$ (8)
where $f(|H|^{2}/\chi^{2})$ being the arbitrary function preserving the Weyl invariance but breaking the $SO(1,4)$ isometry explicitly. In this case, we get the relation between the Jordan frame potential and the frame function by $V_{J}\propto\Omega(H)^{2}$. For the Higgs pole inflation [16], we choose
$f(|H|^{2}/\chi^{2})=m^{2}_{H}\langle\chi^{2}\rangle\,\cdot\frac{|H|^{2}}{\chi^{2}}+\lambda_{H}\langle\chi^{4}\rangle\,\cdot\frac{|H|^{4}}{\chi^{4}}.$ (9)
After the gauge fixing of the Weyl symmetry with $\langle\chi\rangle=\sqrt{6/(1+a)}$, the small violation of the non-compact isometry leads to the Einstein frame potential, $V_{E}=V_{J}/\Omega^{2}=\lambda_{H}|H|^{4}$. Moreover, the Weyl gauge field couples to gravity minimally and it gets mass, $m^{2}_{w}=6ag^{2}_{w}/(1+a)$ [16], after the gauge fixing. The extra couplings due to the Weyl covariant derivatives do not change the inflationary predictions of the Higgs pole inflation much for $|a|\ll 1$ [16]. A similar discussion can be applied to the case of the PQ pole inflation.
@@ 3.2 Coleman-Weinberg potential for the Higgs inflaton
where $\alpha=\{Z,W,t,h,G\}$ for $Z,W$ gauge bosons, top quark, Higgs and Goldstones, respectively, with $N_{\alpha}=\{3,6,-12,1,3\}$, and $C_{\alpha}=\frac{3}{2}$ for fermions or scalars and $C_{\alpha}=\frac{5}{6}$ for gauge bosons. For large Higgs field values during inflation, we choose the renormalization scale to $\mu=c_{*}h/\sqrt{\Omega}$ with $c_{*}$ being the optimal value of order one in order to keep the logarithms minimized in the perturbative expansion of the Coleman-Weinberg potential [19] as the Higgs field value changes. One could choose another renormalization scale, for instance, a field-independent constant value. In this case, the CW potential is field-dependent and dominated by $\ln(h/\sqrt{\Omega})$, while $\lambda(\mu)$ in the tree-level potential is constant, leading to the same results as in the case with $\mu=c_{*}h/\sqrt{\Omega}$. So, the results do not depend on the choice of the renormalization scale [20]. However, in the case with a constant $\mu$, there can be large logarithms in the CW potential as the Higgs field varies, so the scale independence would not be manifest at a finite order of the loop expansion [21].
Following the procedure described in the Appendix A, we find that the renormalized effective potential during inflation is expanded for $h^{2}<6$ as
@@
where the running quartic coupling $\lambda_{H}$ becomes field-dependent due to $\mu=c_{*}h/\sqrt{\Omega}$ and we introduced the sextet coupling $\lambda_{6}$ for the Higgs after the renormalization of the CW potential [^1] See the details for the renormalization of the CW potential at one-loops in the Appendix A.. Henceforth, choosing the renormalized sextet coupling and the higher order self-couplings for the Higgs to small values by the renormalization conditions such that they are sub-dominant for inflation, we focus on the effects of the running Higgs quartic coupling. We note that the Higgs and Goldstone contributions to the CW potential can be suppressed, relative to $W,Z$ boson and top quark contributions, but they are important for deriving the RG equations as shown in the Appendix A.
@@
The running quartic coupling can be positive up to $\sim 10^{10}\,{\rm GeV}$ around the inflation scale even within the SM within the certainty of the pole mass measurement for top quark [24]. It is also possible to make the beta function for the Higgs quartic coupling with new physics contributions to be small around the inflation scale if the new physics scale appears below the instability scale of the SM vacuum [13].
Figure 1: (Left) The full two-loop running Higgs quartic coupling $\lambda_{H}$ for the SM in black dashed line and for the SM plus the singlet scalar in blue solid line. (Right) The running Higgs $\lambda_{H}$ and its beta function in blue and black lines, respectively. The beta function remains negative all the way to $\mu\sim M_{P}$.
Figure 2: The same as in Fig. 1, except for relatively larger couplings for the singlet scalar. The beta function becomes a positive value near $\mu\sim M_{P}$.
In Fig. 1, we depict the running Higgs quartic coupling from the two-loop renormalization group equations [19, 22] in blue lines, after the tree-level shift and loop corrections due to the singlet scalar with mass $m_{S}=10^{9}\,{\rm GeV}$ are included on both sides and the beta function for the Higgs quartic coupling is shown in black line in the right plot. For comparison, the running Higgs quartic coupling in the SM is shown in dashed black line in the left plot. The Higgs quartic coupling is positive and small while the beta function remains negative and small toward $\mu\sim M_{P}$. Here, we chose the singlet quartic coupling to $\lambda_{S}=0.05$ and $\lambda_{HS}=0.02739$ at the singlet threshold.
In Fig. 2, we also show the case for $\lambda_{S}=0.3$ and $\lambda_{HS}=0.06124$ at the singlet threshold with $m_{S}=10^{9}\,{\rm GeV}$. In this case, the Higgs quartic coupling gets positive and small near the Planck scale and the beta function for the Higgs quartic coupling also turns positive and small.
Assuming that new physics contributions to the running of the Higgs quartic couplings exists in terms of the tree-level shift in eq. (31) and the loop corrections in the Coleman-Weinberg potential with eq. (30), we parametrize the running Higgs quartic coupling as being running up to two-loops near the inflation scale $\mu_{I}$, as follows,
@@ 3.3 The case for the PQ inflaton
with the field-dependent quartic coupling $\lambda_{\Phi}(\rho)$ and the sextet coupling ${\bar{\lambda}}_{6}$ introduced after the renormalization of the CW potential. Similarly to the Higgs pole inflation, we choose the renormalization conditions for the renormalized sextet coupling and the higher order self-couplings for the PQ field by setting them to small values at the inflation scale to be sub-dominant during inflation, so we focus only on the running quartic coupling for the PQ field up to two-loops as in eq. (32) in the previous section. But, the beta function coefficients, $b_{1},b_{2}$, in eq. (32), are model-dependent, and they can be different from those in the Higgs pole inflation.
For the PQ pole inflation, the beta function coefficients of the Higgs quartic coupling also depend on $y_{Q},y_{N}.,\lambda_{H\Phi}$ in KSVZ models and $y_{N},\kappa,\lambda_{1\Phi},\lambda_{2\Phi}$ in DFSZ models, in contrast to the Higgs pole inflation where a singlet scalar field can be added to improve the vacuum instability. Thus, in this case, a small running quartic coupling for the PQ field is maintained in the presence of small couplings of the PQ fields, namely, as far as $y_{Q},y_{N}\lesssim 10^{-3}$ and $\lambda_{H\Phi},\kappa,\lambda_{1\Phi},\lambda_{2\Phi}\lesssim 10^{-5}$ [14, 15], and the beta function can be chosen to be small and of either signs for appropriate choices of new couplings. Therefore, we don’t pursue a concrete analysis of the running quartic coupling for the PQ field.
@@ 4.2 Inflationary observables
$=$ $3\int^{\phi_{*}}_{\phi_{e}}\bigg(\frac{4\sqrt{6}}{\sinh\big(\frac{2\phi}{\sqrt{6}}\big)}+\frac{3\sqrt{6}b_{1}+6b_{2}(\phi-\phi_{e})}{6\lambda_{H,e}+\sqrt{6}b_{1}(\phi-\phi_{e})+b_{2}(\phi-\phi_{e})^{2}}\bigg)^{-1}d\phi.$
@@
$N$ $\simeq$ $\frac{3}{4\sqrt{6}}\int^{\phi_{*}}_{\phi_{e}}d\phi\,\sinh\Big(\frac{2\phi}{\sqrt{6}}\Big)\bigg(1-\frac{3b_{1}+\sqrt{6}b_{2}(\phi-\phi_{e})}{24\lambda_{H,*}}\,\sinh\Big(\frac{2\phi}{\sqrt{6}}\Big)\bigg)$ (55)
@@
The observed spectral index is given by $n_{s}=0.9649\pm 0.0044$ from Planck [2], $n_{s}=0.9709\pm 0.0038$ from Planck++ACT [3], and $n_{s}=0.9743\pm 0.0034$ from Planck++ACT++LB [3]. On the other hand, the bound on the tensor-to-scalar ratio from the combined Planck and Keck data is given by $r<0.036$ at $95\%$ CL [6].
@@
Figure 3: Parameter space for $b_{1}/\lambda_{H_{*}}$ vs $b_{2}/\lambda_{H_{*}}$ being consistent with Planck++ACT++LB data [3]. The range of the spectral index within $1\sigma$ and $2\sigma$ errors are shown in yellow and green. We took $N=50$ on left and $N=60$ on right. The gray region is not consistent with perturbativity and the purple region shows the two-loop dominance.
@@
Figure 5: The spectral index vs the tensor-to-scalar ratio. The range of the spectral index allowed by Planck++ACT++LB data [3] are shown in yellow and green colors within $1\sigma$ and $2\sigma$ errors, respectively. We chose $b_{1}/\lambda_{H_{*}}=0.02(-0.01)$ and $b_{2}/\lambda_{H_{*}}=0.001(0.005)$ on left and right, respectively.
In Fig. 3, we depict the parameter space for the one-loop and two-loop beta function coefficients, namely, $b_{1}/\lambda_{H_{*}}$ vs $b_{2}/\lambda_{H_{*}}$, that are both positive and are consistent with Planck++ACT++LB within $1\sigma$ and $2\sigma$ errors [3]. We took $N=50,60$ in the left and right plots, respectively. The gray region is not consistent with perturbativity conditions for loop corrections while the purple region is where the two-loop corrections are larger than the one-loop corrections.
@@
In Fig. 5, we also show the predictions for the spectral index vs the tensor-to-scalar ratio in our model. We chose $b_{1}/\lambda_{H_{*}}=0.02$ and $b_{2}/\lambda_{H_{*}}=0.001$ in the left plot while $b_{1}/\lambda_{H_{*}}=-0.01$ and $b_{2}/\lambda_{H_{*}}=0.005$ in the right plot. For N=5060N=50-60, the inflationary predictions can be read along the black line bounded by two black bullets. In these examples, for N=5060N=50-60, the spectral index and the tensor-to-scalar ratio vary by $n_{s}=0.967-0.974$ and r=0.00500.0068r=0.0050-0.0068 on left and $n_{s}=0.966-0.973$ and r=0.00400.0057r=0.0040-0.0057 on right. As a result, the running quartic coupling at the horizon exit in eq. (61) varies by $\lambda_{H,*}=(1.7-2.3)\times 10^{-11}$ on left and $\lambda_{H,*}=(1.4-2.0)\times 10^{-11}$ on right. The results are in contrast with the case at tree level for $b_{1}=b_{2}=0$: $n_{s}=0.960-0.967$ and r=0.00330.0048r=0.0033-0.0048 for N=5060N=50-60, so the predictions for the spectral index are in a tension with Planck++ACT++LB at about $2-4\sigma$ level.
@@ 5 Conclusions
We presented the Coleman-Weinberg potential for the inflaton in the pole inflation scenarios through the running quartic coupling for the inflaton. The loop corrections stem from the SM particles and extra singlet scalar fields in the Higgs pole inflation, giving rise to the inflaton-dependent power corrections to the Higgs quartic coupling during inflation, although the non-inflaton fields are much heavier than the inflaton. Similar power corrections are obtained for the PQ pole inflation, due to the couplings of the PQ field in the realizations of the PQ symmetry: the Yukawa couplings of the heavy vector-like quark and the right-handed neutrinos and the portal coupling to the SM Higgs doublet in KSVZ models, and the Yukawa couplings of the right-handed neutrinos and the portal couplings to two Higgs doublets in DSFZ models. Thus, the concrete realizations of the pole inflation such as the beta function coefficients of the quartic coupling, $b_{1}$ and $b_{2}$, are model-dependent, but we can parametrize the running quartic coupling by the same function of the inflaton in a model-independent way, because the effective masses of the spectators fields in both models depend on the pole of the inflaton in the same way.
We showed that the loop corrections can make the inflationary predictions in the pole inflation deviating from the tree-level results, in favor of the ACT results. For a positive one-loop beta function for the inflaton quartic coupling (namely, $b_{1}>0$), a sub-dominant contribution from the two-loop corrections can be accommodated. On the other hand, if the one-loop beta function for the inflaton coupling is negative (namely, $b_{1}<0$), we need sizable contributions from two-loops that are larger than the one-loop corrections to explain the ACT results. Whether $b_{1}>0$ or $b_{1}<0$, we found that the tensor-to-scalar ratio remains small enough to be compatible with the bound from Planck and Keck, although it gets slightly larger due to the loop corrections.